OpenStudy (anonymous):

What is the solution of the system of equations? (Matrix) 2x+2y+3z=-6 3x+5y+4z=3 2x+3y+4z=-10

OpenStudy (anonymous):

@satellite73 Please assist me with this!

OpenStudy (habbster):

via matrix operation?

OpenStudy (habbster):

step by step?

OpenStudy (anonymous):

yes to both. :D

OpenStudy (anonymous):

Well you can solve it the easiest way which I do believe is matrix (which I suck at)

OpenStudy (habbster):

wait ill try

OpenStudy (anonymous):

k :D

OpenStudy (habbster):

do you need the elementary row operation?

OpenStudy (anonymous):

Not sure exactly what you mean.

OpenStudy (habbster):

the one you need to form it into echelon form

OpenStudy (anonymous):

You don't need matrix

OpenStudy (anonymous):

Elimination method is the best.

OpenStudy (anonymous):

How would I use that on this?

OpenStudy (habbster):

@Azteck i think he is required to use matrix

OpenStudy (anonymous):

No, just to solve the equation

OpenStudy (habbster):

then just use elimination

OpenStudy (anonymous):

But what would be the best way to do that?

OpenStudy (anonymous):

\[2x+2y+3z=-6[1]\] \[3x+5y+4z=3[2]\] \[2x+3y+4z=-10[3]\] \[[3]-[1]\] \[(2x+3y+4z)-(2x+2y+3z)=(-10)-(-6)\] \[y+z=-10+6\] \[y+z=-4[4]\] \[[3]\times 3\] \[3(2x+3y+4z)=3(−10)\] \[6x+9y+12z=-30[5]\] \[[2]\times 2\] \[2(3x+5y+4z)=2(3)\] \[6x+10y+8z=6[6]\] \[[6]-[5]\] \[(6x+10y+8z)-(6x+9y+12z)=(6)-(-30)\] \[y-4z=6+30\] \[y-4z=36[7]\] \[[4]-[7]\] \[(y+z)-(y-4z)=(-4)-(36)\] \[5z=-40\]

OpenStudy (anonymous):

\[z=-8\]

OpenStudy (anonymous):

To find the y-value, substitute z into one of the equations that contain two variables (ie. [4] or [7])

OpenStudy (anonymous):

Once you find the y-value, you can find x by subbing the two values you just found into one of the two equations in the first two lines above. (ie. [1] or [2] or [5] or [6])

OpenStudy (anonymous):

So basically I'm spposed to have: 2x+2y+3z=-6 become 2x+2y+3(-8)=-6 and then 2x+2y+-5=-6?

OpenStudy (anonymous):

Is this the quickest way to do it? ^^;;

OpenStudy (anonymous):

No... There were two equations above that had two variables.

OpenStudy (anonymous):

Find equation [4] and equation [7] above.

OpenStudy (anonymous):

Use either one of them to find y.

OpenStudy (anonymous):

ohh okay

OpenStudy (anonymous):

So somethng like : y+z=-10+6 becomes y+z=-4 so -8-4=y right?

OpenStudy (anonymous):

I suggest you copy down all the working out I gave and then read it again in your exercise book to better understand it in hard copy what I did to get z.

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

No.

OpenStudy (anonymous):

\[y+(-8)=-4\]

OpenStudy (anonymous):

\[y-8=-4\]

OpenStudy (anonymous):

When you move -8 to the RHS(Right Hand Side), wouldn't the negative sign change?

OpenStudy (anonymous):

@xKingx

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Then why didn't you change it? Need to be careful with all your basics if you want to solve these types of questions. one mistake can cost you full marks for the question.

OpenStudy (anonymous):

right srry, guess I'm just tired. x.x so y=4?

OpenStudy (anonymous):

Correct.

OpenStudy (anonymous):

Now you can use one fo the 4 equations above that had 3 variables to find x.

OpenStudy (anonymous):

use one of*

OpenStudy (anonymous):

So for 3(2x+3y+4z)=3(-10). We divide 3 on both sides and get (2x+3(4)+4(-8)=(-10) or rather: 2x+12-32=(-10) so 2x+-20=-10. On the right track so far? (Again I'm dead tired so brains a little slow atm)

OpenStudy (anonymous):

Use an equation with a number beside it (ie. [?])

OpenStudy (anonymous):

If there are brackets around a number next to an equation and the equation has three variables, use the equation. Why are you taking equations that are in the middle of being multiplied?

OpenStudy (anonymous):

If you were tired, you would be taking advantage of that working out and finding the values quickly in order to go to bed.

OpenStudy (anonymous):

Tired doesn't mean that you're high or drunk or anything like that.

OpenStudy (anonymous):

Lol...? Depends on how tired you are and how many random questions of math you've had to solve for hours straight. In any case: 2x+3y+4z=-10[3]. 2x+3(4)+4(-8)=-10[3]. 2x+12-32=-10[3]. 2x+-20=-10[3]. Like that?

OpenStudy (anonymous):

You smashed it.

OpenStudy (anonymous):

Awesome work.

OpenStudy (anonymous):

kewl. so x=-5 right?

OpenStudy (anonymous):

Remember the sign changes when you move term to the other side of the equal sign.

OpenStudy (anonymous):

when you move a term*

OpenStudy (anonymous):

oh yeah! x.x just 5 then

OpenStudy (anonymous):

Good job.

OpenStudy (anonymous):

Thanks so much, ur awesome lolz.

OpenStudy (anonymous):

No worries.